### v1

# Fact-Checking the Output of Large Language Models via Token-Level Uncertainty Quantification

Abstract
Large language models (LLMs) are notorious for hallucinating, i.e., producing erroneous claims in their output. Such hallucinations can be dangerous, as occasional factual inaccuracies in the generated text might be obscured by the rest of the output being generally factual, making it extremely hard for the users to spot them. Current services that leverage LLMs usually do not provide any means for detecting unreliable generations. Here, we aim to bridge this gap. In particular, we propose a novel fact-checking and hallucination detection pipeline based on token-level uncertainty quantification. Uncertainty scores leverage information encapsulated in the output of a neural network or its layers to detect unreliable predictions, and we show that they can be used to fact-check the atomic claims in the LLM output. Moreover, we present a novel token-level uncertainty quantification method that removes the impact of uncertainty about what claim to generate on the current step and what surface form to use. Our method Claim Conditioned Probability (CCP) measures only the uncertainty of particular claim value expressed by the model. Experiments on the task of biography generation demonstrate strong improvements for CCP compared to the baselines for six different LLMs and three languages. Human evaluation reveals that the fact-checking pipeline based on uncertainty quantification is competitive with a fact-checking tool that leverages external knowledge.
Fact-Checking the Output of Large Language Models
via Token-Level Uncertainty Quantification
Ekaterina Fadeeva3,4 $\diamondsuit$  Aleksandr Rubashevskii1,3 $\diamondsuit$  Artem Shelmanov1 $\diamondsuit$ Sergey Petrakov3 Haonan Li1 Hamdy Mubarak7 Evgenii Tsymbalov8 Gleb Kuzmin2,5 Alexander Panchenko2,3  Timothy Baldwin1,6  Preslav Nakov1  Maxim Panov1 1MBZUAI 2AIRI 3Center for Artificial Intelligence Technology  4HSE University 5FRC CSC RAS 6The University of Melbourne 7QCRI 8Independent Researcher {ekaterina.fadeeva, sergey.petrakov}@skol.tech    {kuzmin, panchenko}@airi.net {aleksandr.rubashevskii, artem.shelmanov, haonan.li}@mbzuai.ac.ae {maxim.panov, timothy.baldwin, preslav.nakov}@mbzuai.ac.aehmubarak@hbku.edu.qa
1 Introduction
footnotetext: $\diamondsuit$ Equal contribution
Large language models (LLMs) have become a ubiquitous and versatile tool for addressing a variety of natural language processing (NLP) tasks. People use these models for tasks including information search Sun et al. (2023b), to ask medical questions Thirunavukarasu et al. (2023), or to generate new content Sun et al. (2023a). There has been notable shift in user behavior, indicating an increasing reliance on and trust in LLMs as primary information sources, often over traditional channels. However, a significant challenge that arises with the spread of these models is their tendency to produce “hallucinations”, i.e., factually incorrect generations featuring misleading information Bang et al. (2023); Dale et al. (2023). This is a side-effect of the way modern LLMs are designed and trained Kalai and Vempala (2023).
LLM hallucinations are a major concern because the deceptive content at the the surface-level can be very coherence and persuasive. Common examples include the creation of fictitious biographies or the assertion of unfounded claims. The danger is that a few occasional false claims might be easily obscured by a large number of factual statements, making it extremely hard for people to spot them. As hallucinations in LLM outputs are hard to eliminate completely, users of such systems could be informed via highlighting some potential caveats in the text, and this is where our approach can help.
Fact-checking is a research direction that addresses this problem. It is usually approached using complex systems that leverage external knowledge sources Guo et al. (2022); Nakov et al. (2021); Wadden et al. (2020). This introduces problems related to the incomplete nature of such sources and notable overhead in terms of storing the knowledge. We argue that information about whether a generation is a hallucination is encapsulated in the model output itself, and can be extracted using uncertainty quantification (UQ) techniques Gal et al. (2016). This avoids implementing complex and expensive fact-checking systems that require additional computational overhead and rely on external resources.
Prior work has mainly focused on quantification of uncertainty for the whole generated text and been mostly limited to tasks such as machine translation Malinin and Gales (2020), question answering Kuhn et al. (2022), and text summarization van der Poel et al. (2022). However, the need for UQ of only a portion of the generations substantially complicates the problem. We approach it by leveraging token-level uncertainty scores and aggregating them into claim-level scores. Moreover, we introduce a new token-level uncertainty score, namely claim-conditioned probability (CCP), which demonstrates confident improvements over several baselines for six LLMs and three languages.
To the best of our knowledge, there is no previous work that has investigated the quality of claim-level UQ techniques for LLM generation. Therefore, for this purpose, we construct a novel benchmark based on fact-checking of biographies of individuals generated using a range of LLMs. Note that different LLMs produce different outputs, which generally have higher variability than, e.g., outputs in such tasks as machine translation or question answering. Therefore, we compare the predictions and uncertainty scores to the results of an automatic external fact-checking system FactScore Min et al. (2023). Human evaluation verifies that our constructed benchmark based on FactScore can adequately evaluate the performance of the uncertainty scores.
Our contributions are as follows:
  • We propose a novel framework for fact-checking LLM generations using token-level uncertainty quantification. We provide a procedure for efficiently estimating the uncertainty of atomic claims generated by a white-box model and highlighting potentially deceptive fragments by mapping them back to the original response.
  • We propose a novel method for token-level uncertainty quantification that outperforms baselines and can be used as a plug-in in a fact-checking pipeline.
  • We design a novel approach to evaluation of token-level UQ methods for white-box LLMs based on fact-checking, which can be applied to other white-box LLMs.
  • We provide an empirical and ablation analysis of method for fact-checking of LLM generations, and find that the uncertainty scores we produce can help to spot claims with factual errors for six LLMs over three languages: English, Chinese and Arabic.
  • All code and data will be made available upon acceptance.
Maximum Probability
Refer to caption Tell me a bio of Madonna. Refer to caption Madonna is an American singer, songwriter, actress, and businesswoman. She was born on August 16, 1958, in Bay City, Michigan, and is best known for her provocative image and music . Madonna has sold over 300 million records worldwide, making her one of the best-selling music artists of all time. She has won numerous awards, including 20 Grammy Awards, a Golden Globe Award, and an Emmy Award. Madonna is also known for her philanthropic work and her activism on various social issues.
Claim-Conditioned Probability (ours)
Refer to caption Tell me a bio of Madonna. Refer to caption Madonna is an American singer, songwriter, actress, and businesswoman. She was born on August 16, 1958, in Bay City, Michigan, and is best known for her provocative image and music. Madonna has sold over 300 million records worldwide, making her one of the best-selling music artists of all time. She has won numerous awards, including 20 Grammy Awards , a Golden Globe Award, and an Emmy Award. Madonna is also known for her philanthropic work and her activism on various social issues.
Figure 1: Visual comparison of our proposed Claim-Conditioned Probability method to the Maximum Probability baseline. CCP accurately identifies the incorrectly specified number of awards (in red), whereas Maximum Probability erroneously highlights the claim that is actually correct.
2 Related Work
2.1 Fact-Checking LLM Generations and Detecting Hallucinations
The problem of hallucinations has made fact-checking of LLM outputs a prominent topic in the research community, and resulted in a surge of work. Chern et al. (2023) present Factool – a task and domain agnostic framework for hallucination detection that leverages GPT for claim extraction and verification. Manakul et al. (2023) suggest to sample multiple outputs from black-box LLMs and evaluate how similar the sampled responses are using the external model. Varshney et al. (2023) detect LLM hallucinations by extracting key parts of output using an external model and estimating their uncertainty based on logits. The most uncertain parts are verified using an external knowledge source. Pan et al. (2023) propose to fact-check complex statements by decomposing them into simpler subtasks and generating reasoning programs to verify these statements. Min et al. (2023) present a methodology for evaluating long LLM-generated texts by decomposing them into simple atomic statements and further verifying them against some knowledge source.
In contrast to previous work that leverages external knowledge sources for fact-checking (a database or another LLM), our work is the first to investigate token-level UQ methods for this task while having access to the LLM and its outputs only.
2.2 Uncertainty Quantification of LLM Generation
UQ techniques for LLM generation can be classified into five major categories Fadeeva et al. (2023). Information-based methods leverage the probability distribution of generated tokens and usually do not require additional models. In this category, we can include methods such as perplexity Fomicheva et al. (2020), mean token entropy Fomicheva et al. (2020), point-wise mutual information Takayama and Arase (2019), and conditional point-wise mutual information van der Poel et al. (2022).
Another category is based on density estimation of latent model representations. It includes Mahalanobis distance Lee et al. (2018) with various modifications Ren et al. (2023). The disadvantage of such approaches is the need for access to the LLM training data in order to fit external density models, which is infeasible for most LLMs.
Ensembling and Monte Carlo dropout methods are based on the lexical diversity of multiple LLM outputs for a single query Malinin and Gales (2020); Fomicheva et al. (2020). Their main drawback is that they require many predictions, which makes them too computationally and memory intensive for practical purposes.
It has recently been shown that LLMs can reflexively estimate the confidence of their generations simply by asking themselves about the truthfulness of their output Kadavath et al. (2022). This can work better than analyzing the probability distribution for the original prediction, but requires a second pass of inference, feeding the original output as a part of the query.
Finally, there is a group of methods that leverages diversity of meanings that the LLM generates for a given query. This group includes semantic entropy Kuhn et al. (2022) and various scores based on the analysis of the similarity matrix between outputs Lin et al. (2023).
UQ methods can also be classified into white-box and black-box approaches depending on the required access to LLM itself and its outputs. Black-box techniques do not require any other input except the model outputs.
The method in our work can be attributed to the information-based group, and can be applied only to white-box LLMs because it requires access to the probability distribution of the generated tokens. Compared to other techniques, it offers a novel approach to post-processing the probability distribution and is designed to quantify uncertainty in parts of the output, such as atomic claims and individual words.
3 Fact-Checking Pipeline
The fact-checking pipeline (see Figure 5) starts with splitting a generated text into atomic claims, e.g. using a much smaller model fine-tuned for this particular task. For experimental evaluations in this work, we follow the approach of FactScore Min et al. (2023), where splitting is implemented via the ChatGPT API.
Each atomic claim is matched against the sequence of tokens in the original text with corresponding probability distributions. Then we calculate token-level uncertainty scores and aggregate them into the claim-level uncertainty.
Finally, the claim-level uncertainty scores are compared against a threshold obtained on a validation set to determine whether the claim should be highlighted for the end-user as unreliable. Individual tokens can be a part of multiple atomic claims. If the token belongs to a reliable and unreliable claim at the same time, it is not highlighted. An example visualization is presented in Figure 1.
Figure 2: Example of CCP calculation on the Vicuna 13b generation for the word painting.
4 Uncertainty Quantification
In this section, we first provide background on common UQ methods that can be used at the token level, then delve into our Claim-Conditioned Probability (CCP) token-level method, and finally describe how token-level uncertainties are aggregated into a claim-level score.
Autoregressive language models generate text token by token. In this work, we will operate on the level of words and without loss of generality suppose that the autoregressive distribution at each step generates the random word $X_{j}\sim P(\cdot\mid x_{<j})$, where $x_{<j}$ is the text generated before the word at position $j$. We also denote by $x_{j}$ the generated word at position $j$ and by $x_{1:j}=x_{<j}\circ x_{j}$ a text composed of words at positions $1$ to $j$. For example, in the case of greedy generation, $x_{j}=\arg\max_{x}{P(x\mid x_{<j})}$, where $x_{j}$ is the most probable realization of $X_{j}$. Let us also denote by $C$ a set of indices of words corresponding to a particular atomic claim.
4.1 Claim-Level Uncertainty Quantification Baselines
We note that for UQ to be practical, it needs to be fast. Therefore, we do not consider methods such as deep ensembles Lakshminarayanan et al. (2017), due to their significant computational overhead.
Maximum Probability represents a basic approach to UQ, where we simply treat the probability of the most likely generation as a confidence score:
$MP(C)=1-\prod_{j\in C}P(x_{j}\mid x_{<j}).$ (1)
Maximum Entropy of a token in the claim:
$Ent(C)=\max_{j\in C}\mathcal{H}(\cdot\mid x_{<j}),$ (2)
where $\mathcal{H}(\cdot\mid x_{<j})$ is the entropy of the autoregressive distribution of the current token. Preliminary experiments indicated that simply getting the maximum of the token entropies in a claim noticeably outperforms other aggregation techniques like average or minimum. It is also generally a slightly better baseline than perplexity.
P(True), similar to Kadavath et al. (2022), measures the uncertainty of the claim by asking the LLM itself whether the generated claim is true or not, as calculated by the model confidence of the first generated token $y_{1}$ being equal to “True”:
$P(\text{True})=1-P(y_{1}=\text{``True''}).$ (3)
While some work has reported that this technique outperforms other baselines, the big drawback is that one needs to run the original LLM twice.
4.2 Claim-Conditioned Probability
In this subsection, we propose a novel method for token- and claim-level uncertainty quantification.
4.2.1 Motivation and Theoretical Background
When a LLM generates an output, it faces various types of uncertainty reflected in the token distribution of the current generation step (see Figure 4 for an example). We identify three distinct types of uncertainty:
  1. 1.
    Claim type/order uncertainty: What claim to generate on the current step? For example, on the current step, a LLM might hesitate between generating a year of graduation of a person and a field of study. A different order of claims, missing claims, or different types of generated claims do not make produced text less factual. Therefore, when performing fact-checking, we should not take this type of uncertainty into account.
  2. 2.
    Surface form uncertainty: What synonyms or hypernyms to use when generating a claim (e.g. “art” or “painting”)? Different surface forms also do not make the text less factual — they might change the style, but not the underlying meaning of the text. Therefore, this type of uncertainty is also not relevant for fact-checking.
  3. 3.
    Claim uncertainty: What specific piece of information to relay for a particular claim type? For example, an LLM might not be sure what field of study to generate, producing a token distribution with multiple highly-probable variants, such as “painting”, “acting”, “sculpture”. In a similar way, for the year of graduation, a LLM might produce a distribution with various potential years. This uncertainty is relevant for fact-checking, because if the model is not sure about the information it relays, there might be a high chance of a factual mistake.
Two out of the three types of uncertainty are irrelevant for fact-checking and only introduce noise into the final score. We propose a new UQ method that ignores the first two types of uncertainty and focuses only on the third one, namely Claim-Conditioned Probability (CCP):
$CCP(x_{j})=P\bigl(\text{Meaning}(x_{1:j})\mid x_{<j},\text{ClaimType}(x_{1:j})\bigr).$ (4)
Here, $\text{ClaimType}(x_{1:j})$ represents a claim type of the generated sequence $x_{1:j}$ and $\text{Meaning}(x_{1:j})$ is a function that maps $x_{j}$ into its meaning given the previous words in a sentence $x_{<j}$, so that various surface forms with a similar meaning for $x_{j}$ are mapped to a single categorical variable.
Conditional probability can be rewritten using unconditional probabilities:
$P\bigl(\text{Meaning}(x_{1:j})\mid\text{ClaimType}(x_{1:j}),x_{<j}\bigr)$
$=\frac{P(\text{Meaning}(x_{1:j}),\text{ClaimType}(x_{1:j})\mid x_{<j})}{P(\text{ClaimType}(x_{1:j})\mid x_{<j})}.$
Assuming that each meaning in a word distribution can correspond to only a single claim type, the joint probability is the same as the meaning probability $P(\text{Meaning}(x_{1:j}),\text{ClaimType}(x_{1:j})\mid x_{<j})=P(\text{Meaning}(x_{1:j})\mid x_{<j})$.
Meaning probability in turn sums from the probabilities of word alternatives $x^{k}_{j}$ that correspond to the same meaning: $P(\text{Meaning}(x_{1:j})\mid x_{<j})=\sum_{x^{k}_{j}\in M(x_{j})}{}{P(x^{k}_{j}\mid x_{<j})}$, where we say that $x^{k}_{j}\in M(x_{j})$ if $\text{Meaning}(x_{1:j})=\text{Meaning}(x_{<j}\circ x^{k}_{j})$.
In the same way, the probability of a claim type sums from probabilities of all meanings and transitionally from probabilities of words that correspond to the particular claim type: $P\bigl(\text{ClaimType}(x_{1:j})\bigr)=\sum_{x^{l}_{j}\in CT(x_{j})}{}{P(x^{l}_{j}\mid x_{<j})}$, where we denote by $x^{l}_{j}\in CT(x_{j})$ an event such that $\text{ClaimType}(x_{1:j})=\text{ClaimType}(x_{<j}\circ x^{l}_{j})$. Therefore, Equation (4) can be rewritten as follows:
$CCP(x_{j})=\frac{\sum_{x^{k}_{j}\in M(x_{j})}{}{P(x^{k}_{j}\mid x_{<j})}}{\sum_{x^{l}_{j}\in CT(x_{j})}{}{P(x^{l}_{j}\mid x_{<j})}}.$ (5)
The meaning function and the construction of the claim type set of words can be implemented in various ways. We outline one approach in Section˜4.2.2.
Previously-proposed UQ methods have partially accounted for some of the types of uncertainty described above. For example, semantic entropy Kuhn et al. (2023) accounts for uncertainty in semantically-equivalent groups, which helps to alleviate the surface-form uncertainty. However, in our method, we additionally remove the impact of claim-type uncertainty.
4.2.2 Implementation
We implement CCP using NLI at the word level. We compare the original claim and the claim, in which the target word is replaced by its alternatives from the autoregressive distribution.
The distribution $X_{j}$ at the position $j$ is approximated by top-$K$ alternatives $\left\{x^{k}_{j}\right\}_{k=1}^{K}$ with $x^{1}_{j}\equiv x_{j}$. We replace $x_{j}$ with its alternatives $x_{j}^{k}$ and obtain new instances $x_{<j}\circ x^{k}_{j},k=1,\dots,K$. Each new instance is compared against the original prediction $x_{1:j}=x_{<j}\circ x_{j}$ using an NLI model. We define $\texttt{NLI}(x_{j}^{k},x_{j}):=\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$, where $\texttt{NLI}(x_{<j}\circ x^{k}_{j},x_{1:j})$ means application of the NLI model to the sentences $x_{<j}\circ x^{k}_{j}$ and $x_{1:j}$. The outcome of NLI procedure is one of three classes: entail (‘e’), contradict (‘c’) or neutral (‘n’). If the new instance entails the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`e'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has the same meaning with $x_{1:j}$ ($x^{k}_{j}\in M(x_{j})$) and corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$). If the new instance contradicts the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`c'}$, then we consider $x_{<j}\circ x^{k}_{j}$ has a different meaning with $x_{1:j}$ ($x^{k}_{j}\notin M(x_{j})$), but corresponds to the same claim type ($x^{k}_{j}\in CT(x_{j})$). Otherwise, if the new instance is neutral w.r.t. the original prediction $\texttt{NLI}(x_{j}^{k},x_{j})=\textrm{`n'}$, then we consider that $x_{<j}\circ x^{k}_{j}$ does not correspond to the same claim type as $x_{1:j}$ ($x^{k}_{j}\notin CT(x_{j})$). Thus, Equation (5) for CCP can be written as follows:
$CCP_{\textit{word}}(x_{j})=\dfrac{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})=\texttt{`e'}}P(x^{k}_{j}\mid x_{<j})}{\sum_{k:\texttt{NLI}(x^{k}_{j},x_{j})\in\left\{\texttt{`e'},\texttt{`c'}\right\}}P(x^{k}_{j}\mid x_{<j})}.$ (6)
For practical considerations, we consider that CCP for function words is always equal to 1. In our experiments, we base this determination on the stop word list from NLTK Bird and Loper (2004).
We note that most transformer LLMs generate sub-word tokens instead of whole words. To obtain the distribution for whole words, we generate one or multiple tokens using beam search with $K$ beams.
To obtain CCP-based claim-level uncertainty, we simply take the product of CCPs of each word from the claim $C$:
$CCP_{\textit{claim}}(C)=1-\prod\nolimits_{j\in C}CCP_{\textit{word}}(x_{j}).$ (7)
An example of calculating the CCP for a fact is presented in Figure 2. Other detailed examples of CCP calculation are available in Appendix A.
5 Benchmark for Evaluation of Claim-Level UQ Methods
We evaluate claim-level UQ techniques and their ability to spot hallucinations on the task of generating biographies. To make the evaluation as close as possible to the real-world scenario, we allow unrestricted generation of biographies from LLMs, sampling multiple outputs from a given LLM. This complicates the automatic evaluation of the fact-checking pipeline, because obtaining gold standard annotation require that we manually annotate all outputs from each model. Therefore, in addition to manual annotation, we annotate claims in generated texts automatically using the FactScore fact-checking tool Min et al. (2023), which has access to an external knowledge source. Such an approach enables completely automatic evaluation and allows us to scale up our experiments.
We generate LLM responses in English, Chinese, and Arabic to $100$ biography prompts. The typical biography prompt is Give me a biography for <person name> in different languages. The set of people was generated by asking GPT-4 to list the most famous people since 1900. The maximum generation length is set to $256$ tokens. If the last sentence of the generation is unfinished (i.e. does not end with any punctuation), it is discarded. We generate responses for the following LLMs: (for English) Vicuna 13b Zheng et al. (2023), Mistral 7b Jiang et al. (2023), Jais 13b Sengupta et al. (2023), and GPT-3.5-turbo Ouyang et al. (2022); (for Chinese) Yi 6B Yi (2023); (for Arabic) Jais 13b and GPT-4.
Model Mistral 7b Vicuna 13b Jais 13b GPT-3.5-turbo
CCP (ours) 0.66 $\pm$ 0.03 0.66 $\pm$ 0.04 0.71 $\pm$ 0.05 0.60 $\pm$ 0.04
Maximum Prob. 0.59 $\pm$ 0.03 0.60 $\pm$ 0.05 0.64 $\pm$ 0.05 0.54 $\pm$ 0.05
Token Entropy 0.60 $\pm$ 0.02 0.60 $\pm$ 0.06 0.63 $\pm$ 0.06 0.53 $\pm$ 0.03
P(True) 0.53 $\pm$ 0.02 0.61 $\pm$ 0.03 0.55 $\pm$ 0.05 0.53 $\pm$ 0.04
Table 1: ROC-AUC of claim-level UQ methods with FactScore labels as the ground truth (English).
Figure 3: ROC-AUC of claim-level UQ methods based on FactScore labels, aggregated into bins when considering only facts from the first 22, 55, and all sentences (English).
We then decompose the generated text into atomic claims using ChatGPT. For each atomic statement, we map all its words back to generated text to access the corresponding token logits. Not all claims perfectly match to the original response. For example, for Vicuna 13b around $5\%$ of all claims do not successfully match because ChatGPT abstained to respond or outputted words not present in the original. We consider only successfully matched claims.
For English, we are able to perform annotation completely automatically: each atomic claim is classified by FactScore as supported or not supported. The underlying FactScore model is “retrieval+ChatGPT” with a dump of Wikipedia articles as an external knowledge source. We also manually annotated 100 English biographies produced by the Vicuna 13b model, 183 claims from Arabic biographies, and 1,603 claims in Chinese. Each of the statements was checked by two annotators with access to the corresponding Wikipedia article. The final label is set to ‘supported’ only if both annotators label it as so.
The statistics of the resulting datasets with automatically-labeled claims in English are presented in Table 4, and the statistics of manually annotated datasets is presented in Table˜5. As we can see, the majority of claims in the model output are correct, with 6-29% of hallucinations.
The automatic dataset construction process using FactScore is illustrated in Figure 6 in the Appendix.
6 Experiments
6.1 Experimental Setup
Fact-checking of atomic claims is framed as a binary classification task, where uncertainty scores serve as predictors of non-factuality, and FactScore or human labels serve as ground truth. The evaluation metric is ROC-AUC.
For the CCP method, NLI scores are calculated using the DeBERTa-large model He et al. (2021) fine-tuned for this task. The number of alternatives used in CCP is $K=10$, except for GPT-3.5-turbo and GPT-4, as the OpenAI API does not allow us to retrieve more than $5$ alternatives from the token distribution. Details on expenses are presented in Appendix H.
6.2 Results for English on the FactScore Annotation
The main results of the experiments with FactScore labels for English are presented in Table 1. The proposed CCP method outperforms all other UQ techniques for each of the considered LLMs. The best overall result is obtained for Jais 13b, where CCP outperforms the closest competitor by 0.07 ROC-AUC.
We further analyze the performance by plotting it as a function of the number of considered sentences from the beginning of the generated text (see Figure 3). We note that the quality of each method decreases as the number of considered sentences increases. This may be due to the fact that the model tends to start the response with easy-to-know and hence reliable claims and as it generates more text, it has to produce more complex and less reliable statements, which is illustrated for the Vicuna 13b model in Figure 8 in the Appendix. For the majority of cases, CCP outperforms other methods except when we consider only the first two and the first five sentences generated by GPT-3.5-turbo.
6.3 Multilingual Results on Manual Annotation
Multilingual results based on manual annotation are presented in Tables˜2 and 3 and Figure˜9 in the Appendix.
Using manual annotation for English, we can also evaluate the performance of FactScore itself. The accuracy of automatic annotation is $77.2\%$ and ROC-AUC is $0.72$. A detailed analysis of FactScore mistakes is presented in Appendix F. For human annotation, the performance of all UQ methods appears to be even slightly higher than for the labels obtained using FactScore (Table 2) on the same set. Moreover, CCP outperforms FactScore itself by 0.06 ROC AUC. These results demonstrate that in the task of detecting LLM hallucinations, UQ techniques can be a strong alternative to fact-checking tools with an external knowledge source.
For Chinese and Arabic, CCP also outperforms the baselines. For the Chinese Yi 6b model, from Figure 9, we can see that the gap between CCP and the baselines is especially big for the several first claims. When considering more claims in the LLM output, CCP still clearly outperforms the maximum probability baseline, but the P(True) baseline substantially reduces the gap. For Jais and Arabic, CCP outperforms the closest competitor by 0.05 ROC AUC. On GPT-4 generations for Arabic, the metrics for all methods are low. We explain this observation by a small ratio of non-factual claims in the GPT-4 output. Compared to the other models, GPT-4 hallucinates much less.
Method Ground-Truth Human FactScore
annotation annotation
CCP (ours) 0.78 0.74
Maximum Prob. 0.67 0.65
Token Entropy 0.69 0.65
P(True) 0.68 0.65
FactScore 0.72
Table 2: ROC-AUC of claim-level UQ methods with human annotation and FactScore annotation as the ground truth (English, Vicuna 13b model).
Model
Yi 6b,
Chinese
Jais 13b,
Arabic
GPT-4,
Arabic
CCP (ours) 0.64 $\pm$ 0.03 0.66 $\pm$ 0.02 0.56 $\pm$ 0.05
Maximum Prob. 0.52 $\pm$ 0.03 0.59 $\pm$ 0.02 0.55 $\pm$ 0.08
Token Entropy 0.57 $\pm$ 0.05 0.61 $\pm$ 0.02 0.48 $\pm$ 0.06
P(True) 0.63 $\pm$ 0.04 0.61 $\pm$ 0.02 0.50 $\pm$ 0.06
Table 3: ROC-AUC of claim-level UQ methods with manual annotation as the ground truth.
6.4 Ablation Studies
In this section, we analyze the influence of various CPP components on English biographies annotated with FactScore (Tables 68). Details of the experimental setup for each ablation study are presented in Appendix D.
(1) Aggregation of $CCP_{\textit{word}}$ for obtaining $CCP_{\textit{claim}}$. Besides the product of probabilities, we also tried the normalized product, minimum and average probability. All these approaches perform slightly worse than the product (see Table 6).
(2) NLI context. We analyze what context is sufficient for NLI in CCP (Table 7). In addition to the standard variant in CCP, where we keep the claim that precedes the word in question, we experiment with a single target word without context and the whole sentence that precedes the target word. All variants demonstrate lower performance. No context leads to a drop in ROC-AUC of 0.02, and longer contexts by more than 0.07.
(3) Functional words handling. The results in Figure 8 show that excluding functional words in CCP helps to improve the performance by 0.03 ROC-AUC. This approach also slightly improves the maximum probability baseline, but its performance is still much lower.
(4) The number of alternatives $K$. Taking $K=5$ alternatives instead of $K=10$ in CCP decreases ROC-AUC by 0.02. Figure 7 also demonstrates that further decreasing $K$ reduces the performance even more. When increasing $K$, the performance plateaus at $K=8$.
6.5 Qualitative Analysis
In qualitative analysis of uncertainty scores for various generations and models, we note that the maximal probability baseline produces a lot more false positives than CCP. This happens because CCP ignores some types of uncertainty, focusing only on the claim uncertainty. In some cases, CCP also finds false claims overlooked by other methods because ignoring certain types of uncertainty also allows us to reduce the cut-off threshold used to mark claims. An example where we compare CCP with maximal probability is presented in Figure 12, and more examples can be found in Appendix G.
7 Conclusion
We presented a novel approach to fact-checking and hallucination detection based on token-level uncertainty quantification. According to human evaluation, our method is competitive with FactScore, a fact-checking tool that leverages an external knowledge source: we achieve similar or better results with access to only LLM outputs.
Our proposed token-level UQ method, claim conditioned probability, outperforms a number of baselines. In this method, we post-process the word distribution to mitigate the impact of uncertainty related to the variability of surface forms and uncertainty about what claim type to generate on the current step. In the constructed benchmark, where we detect hallucinations in biographies, CCP outperforms other methods for six LLMs, including GPT-3.5-turbo and GPT-4, and three languages.
Limitations
We strove to conduct our research according to the best established experimental and methodological practices, yet we identify several potential limitations of this work below.
First, at the core of the approach lies a text entailment classifier. As it was originally pre-trained for a slightly different use-case, more careful analysis of its performance on diverse domains and genres should be carried out.
Second, the current implementation of the method makes use of OpenAI’s GPT models for text segmentation and extraction of atomic facts, similarly to FactScore, which may not be practical in real applications. Replacing these components with cheaper open models should be feasible in principle, but is left for future work.
Third, part of our experimental results rely on a human evaluation, which may be subjective. We tried to mitigate this by creating detailed instructions for the annotators, yet a larger scale study with larger overlap may be desirable.
Fourth, our uncertainty quantification is based on tokens, while taking into account also larger units as noun or verb phrases as basic units of analysis may be better motivated linguistically.
Finally, our approach only detects potentially spurious generations. A prominent direction for future research is to modify the generation of an LLM to exclude such spans, while ensuring fluency of a generated text: a simple removal of uncertain claims may result in incoherent sentences.
Ethical Considerations
We would like to warn that our method for fact-checking is based on uncertainty quantification, and thus it is not bullet-proof, as it only reflects the internal belief of the LLM. Thus, it is of limited utility when it comes to claims that are beyond the time cutoff of the model. Moreover, the LLM might be trained on factually false data, which is beyond our control. It could be also tricked by variations in the prompt.
We also warn that our solution does not eliminate hallucinations; instead, it can be used to highlight risky parts of a text, for humans to take into account. Our intended use is towards raising awareness and promoting human-machine collaboration.
We warn that our method can be misused to unfairly moderate content. Thus, we ask researchers and potential users to exercise caution.
Appendix A Example of CCP Calculation
Figure 4 shows the example of model generation and NLI classes on the sentence from dataset sample. In this example, CCP estimates the Claim Uncertainty, not taking into account Claim order and Surface form uncertainties, in result producing more reasonable uncertainty estimate than Maximum Probability does.
Figure 4: Example part of the bibliography generated by Vicuna 13b model. The words from greedy-generated sentence are presented sequentially on the top, each non-functional word is supplemented with its alternatives and autoregressive generation probabilities. Words with probability less than 0.1% are omitted.
Green-colored words indicate entailment to the greedy generated word, red color indicates contradiction, and yellow color indicates neutral NLI class.
On the last position, CCP successfully distinguishes Norway from other year-related words, and does not consider its probability in the final formula.
Appendix B Fact-Checking Pipeline
Figure 5: Fact-Checking pipeline based on token-level uncertainty quantification.
Appendix C Benchmark Construction Details
C.1 Benchmark Construction Pipeline
Figure 5 presents the suggested general preparation pipeline of a factuality dataset for an arbitrary model and Fact-Checking benchmark. This pipeline was used to generate the biography dataset. Figure 6 presents the detailed pipeline of the biography dataset preparation, based on the general schema.
The prompt for the matching model to map atomic claims to the initial text is as follows: “Given the fact ‘‘fact’’, identify the corresponding words in the original sentence ‘‘sent’’ that help derive this fact. Please list all words that are related to the fact, in the order they appear in the original sentence, each word separated by comma.”. The prompt for the model that partitions the initial generated text into individual atomic claims, as well as the prompt for the classification model, are taken similar to the paper by Min et al. (2023).
C.2 Datasets and Statistics
For Arabic, we generate 100 random biographies from the list of the most visited websites in Arabic Wikipedia to experiment with GPT-4. The used Arabic prompt is the translation of: “Tell me the biography of {person name}”. Then we call GPT-4 to extract atomic claims from the generated Arabic biographies. The used prompt is: “Convert the following biography into Arabic atomic factual claims that can be verified, one claim per line. Biography is: {biography}”. Arabic biographies and claims are translated into English using Google Translate. It’s worth mentioning that almost one-third of the names in the list of person names are foreign, e.g. Donald Trump, Messi, Isaac Newton, Pope Benedict XVI, etc. This can be useful to study the factuality of claims about the same persons in different Wikipedia languages. On average, GPT-4 generates 20 claims from each biography, and random two claims from each biography are verified manually (total = 200 claims).
For Jais 13b experiments, we use the same prompts used in GPT-4. We notice that the biographies generated by Jais 13b are much shorter than the ones generated by GPT-4 (almost half length). We use GPT-4 to extract claims from the generated biographies. On average, biographies generated by Jais 13b have nine claims. Out of the 100 person names, Jais 13b generates empty biographies for seven names with response messages like: “I am sorry! I cannot provide information about {name}”, or “What do you want to know exactly?”. Two random claims from each biography are verified manually (total = 186 claims).
For Chinese, we first prompt ChatGPT to generate a list of 100 famous people. Then use the same way as we have done in Arabic, but change the prompt to Chinese, to generate bios and claims. Matching of Chinese and Arabic claims is performed by aligning the matched symbols to the biography generation greedily. As FactScore pipeline only works with English claims, both Chinese and Arabic datasets were annotated manually. Only claims containing factual information were manually annotated. The statistics for this dataset are presented in Table 5
Model
Number of
claims
Supported
claims
Mistral 7b 3,824 71.1%
Vicuna 13b 3,617 78.2%
Jais 13b 1,407 84.3%
GPT-3.5-turbo 3,875 89.4%
Table 4: The statistics of the datasets generated out of 100 biographies for all tested LLMs for English.
Model
Number of
claims
Supported
claims
Vicuna 13b, English 100 70.1%
Yi 6b, Chinese 1,603 94.0%
Jais 13b, Arabic 186 73.0%
GPT-4, Arabic 200 92.5%
Table 5: The statistics of the datasets generated out of 100 biographies for all tested LLMs.
Figure 6: Example of the algorithm operation of the fact-checking method using uncertainty estimation.
Appendix D Ablation Studies
In this section, we study the influence of different components on the results of CCP method and other baselines.
D.1 Aggregation
Method ROC-AUC
$CCP_{\textit{prod}}$ 0.66 $\pm$ 0.03
$CCP_{\textit{len}}$ 0.65 $\pm$ 0.03
$CCP_{\textit{min}}$ 0.64 $\pm$ 0.03
$CCP_{\textit{mean}}$ 0.65 $\pm$ 0.03
Table 6: ROC-AUC on Vicuna 13b generation, for different normalizations.
Table 6 presents the result of CCP on Vicuna 13b when applying $4$ different kinds of aggregation:
$CCP_{\textit{prod}}(F)$ $=-\prod_{j=1}^{k}CCP_{\textit{word}}(x_{j}),$ (8)
$CCP_{\textit{len}}(F)$ $=-\exp\left(\frac{1}{k}\sum_{j=1}^{k}\log CCP_{\textit{word}}(x_{j})\right),$
$CCP_{\textit{min}}(F)$ $=-\min_{j=1}^{k}CCP_{\textit{word}}(x_{j}),$
$CCP_{\textit{mean}}(F)$ $=-\frac{1}{k}\sum_{j=1}^{k}CCP_{\textit{word}}(x_{j})$
$CCP_{\textit{prod}}$ shows better results and is the one chosen for final $CCP$ formula.
D.2 NLI Context
Method ROC-AUC
$CCP_{\textit{no context}}$ 0.64 $\pm$ 0.03
$CCP_{\textit{sent pref}}$ 0.59 $\pm$ 0.03
$CCP_{\textit{claim pref}}$ 0.66 $\pm$ 0.03
Table 7: ROC-AUC on Vicuna 13b generation, for different contexts to input words with to an NLI model.
Table 7 presents the result of CCP on Vicuna 13b when adding different contexts as the inputs of NLI model:
  1. 1.
    $CCP_{\textit{no context}}$ calculates $NLI$ using these $2$ words as NLI model input;
  2. 2.
    $CCP_{\textit{sent pref}}$ calculates $NLI$ using the prefix in the sentence from model generation up to the current word and its alternative, as input to NLI model;
  3. 3.
    $CCP_{\textit{claim pref}}$, uses all words in the sentence corresponding to the greedy word, which were matched to the current fact and which comes before the greedy word (see example on Figure 2).
The results for sentence prefix provides too wide context to the NLI model, which results in NLI model focusing on the the contexts instead of the greedy words and its alternatives, thus outputting lots of entailment classes. On the other hand, the method without context does not provide enough context to calculate NLI class with more quality. $CCP_{\textit{claim pref}}$ shows better results and is the one chosen for final $CCP$ formula.
D.3 Functional Words Handling
Method ROC-AUC
$CCP_{\textit{confident}}$ 0.66 $\pm$ 0.03
$CCP_{\textit{ignore}}$ 0.63 $\pm$ 0.03
$MP_{\textit{confident}}$ 0.59 $\pm$ 0.05
$MP_{\textit{ignore}}$ 0.60 $\pm$ 0.05
Table 8: ROC-AUC on Vicuna 13b generation, for methods of handling functional words.
Table 8 presents the result of CCP and Maximum Probability on Vicuna 13b when handling functional words differently. The 2 approaches tested are:
  1. 1.
    $CCP_{confident}$ and $MP_{confident}$ are the versions of $CCP$ and $MP$ baselines, which assigns the most confident word-level score of 1.0 to words from NLTK Bird and Loper (2004) stopwords list, which is the same as skipping these words form the list of matched words of length $k$;
  2. 2.
    $CCP_{ignore}$ and $MP_{ignore}$ handles functional words similar to any other words. $MP_{ignore}$ is the baseline used in the main section.
$CCP_{\textit{confident}}$ shows better results and is the one chosen for final $CCP$ formula. In contract, Maximum Probability performs worse if specifically handling functional words.
D.4 The Number of Alternatives
Figure 7 presents the CCP ROC-AUC results on the whole model generation, depending on the number of beams $n$ to run the method with.
Figure 7: ROC-AUC between FactScore classes and claim-level CCP method, dependent on the number of sampled words $n$ to consider.
Appendix E Additional Experimental Results
Here we show additional results related to the performance of our pipeline and CCP method. In the Figure 8 we show the dependence of the percentage of supported claims in the generated LLM response (in this case for Vicuna 13b) as a function of its length. As we can see, the model is more often wrong when the generation length increases, which may be due to the generation of additional facts.
Figure 8: Percentage of supported claims, as a function of the number of sentences to restrict generation to (English, Vicuna 13b).
In the Figure 9, we show the performance quality of our CCP method on Chinese data on the Yi 6b model. We see that, as for other English models, our method outperforms the alternatives over the entire generation length in terms of ROC-AUC.
Figure 9: The comparison of token-level uncertainty quantification methods in terms of ROC-AUC scores, measured for Chinese dataset. The results are split into bins when considering only facts from the first 22, 55, and all sentences.
Appendix F FactScore Annotation System
Here we give examples of the operation of the FactScore automatic markup system, see the Table 9. We hypothesize that the causes of FactScore errors are related to model hallucination (true information is present in the knowledge source, but the model produces incorrect information), lack of context (an excerpt from a Wikipedia article cannot capture all the details), and difficulties with information interpretation (the desired information is present in the knowledge base, but is formulated in different words or in several sentences). All these cases leave room for further improvement of the pipeline. We have also given the results of the proposed CCP method on selected examples. It can be seen that the results of our algorithm are similar to the annotation from FactScore, while CCP does not use external information in calculating the scores.
Cases FS Human CCP Generated Atomic Claim True Information from the Wikipedia article
TN False False 0.819 Marie Stopes died on October 20, 1958. Marie Charlotte Carmichael Stopes (15 October 1880 – 2 October 1958) was a British author…
TN False False 0.999 Planck is best known for his work on the nature of light. His fame as a physicist rests primarily on his role as the originator of quantum theory…
FN False True 1.0 Heisenberg was appointed as the director of the Max Planck Institute. He then became director of the Max Planck Institute for Physics and Astrophysics from 1960 to 1970.
FN False True 0.716 Ray Charles incorporated elements of Latin music into his sound. Charles reached the pinnacle of his success at Atlantic with the release of "What’d I Say", which combined gospel, jazz, blues and Latin music.
FP True False 0.001 Sagan was a prolific writer. Carl Edward Sagan (November 9, 1934 – December 20, 1996) was an American astronomer and science communicator.
FP True False 0.141 Van Gogh was a pastor. Van Gogh prepared for the University of Amsterdam theology entrance examination; he failed the exam …
FP True False 0.185 Hawking showed an early aptitude for science. Although known at school as "Einstein", Hawking was not initially successful academically.
TP True True 0.276 Tiger Woods is a professional golfer. Eldrick Tont "Tiger" Woods (born December 30, 1975) is an American professional golfer.
TP True True 0.015 Miles Davis began playing the trumpet at the age of 13. On his thirteenth birthday his father bought him a new trumpet,[17] and Davis began to play in local bands.
TP True True 0.001 Hitchcock directed "The Birds." Hitchcock’s other notable films include Rope (1948), Strangers on a Train (1951), …, Birds (1963) and Marnie (1964), …
Table 9: Table with examples of the FactScore automatic annotation system and all types of classification outcomes when comparing automatic annotation and manual annotation (confusion matrix elements): True Negative (TN), False Negative (FN), False Positive (FP), True Negative (TN); FS is a FactScore label, Human is a human annotation label. CCP scores are comparable to FactScore annotation and do not require an external source of information.
Appendix G Examples of Results Obtained with CCP and the Maximal Probability Baseline
Maximum Probability
Refer to caption Tell me a bio of Mikhail Gorbachev. Refer to caption Mikhail Gorbachev (born Mikhail Sergeyevich Gorbachev on March 2, 1931, in Stavropol, Russia) is a former Soviet politician and statesman who served as the last President of the Soviet Union from 1989 to 1991. He was also the first president of the Russian Federation from 1991 to 1999.Gorbachev is widely regarded as one of the most influential figures of the 20th century for his role in the collapse of the Soviet Union and the beginning of the end of the Cold War. Gorbachev was born into a peasant family in Stavropol, Russia . …
Claim-Conditioned Probability (ours)
Refer to caption Tell me a bio of Mikhail Gorbachev. Refer to caption Mikhail Gorbachev (born Mikhail Sergeyevich Gorbachev on March 2, 1931, in Stavropol, Russia) is a former Soviet politician and statesman who served as the last President of the Soviet Union from 1989 to 1991 . He was also the first president of the Russian Federation from 1991 to 1999.Gorbachev is widely regarded as one of the most influential figures of the 20th century for his role in the collapse of the Soviet Union and the beginning of the end of the Cold War. Gorbachev was born into a peasant family in Stavropol, Russia. …
Figure 10: Claim-Conditioned Probability vs Maximum Probability baseline.
Maximum Probability
Refer to caption Tell me a bio of Vincent van Gogh. Refer to caption Vincent van Gogh was a Dutch post-Impressionist painter who is widely considered one of the most important figures in the history of art. He was born on March 30, 1853, in Zundert, Netherlands, and died on July 29, 1890, in Auvers-sur-Oise, France. Van Gogh’s early life was marked by a series of unsuccessful careers , including as a teacher , a pastor , and an art dealer . He began to focus on art in his late 20s, and in 1886, he moved to Paris, where he became involved with the avant-garde art scene. In Paris, van Gogh met many other artists, including Claude Monet, Pierre-Auguste Renoir, and Paul Gauguin, who would become his closest friend and mentor. Gauguin’s influence led van Gogh to adopt a more vibrant and expressive style, characterized by bold colors and thick brushstrokes.
Claim-Conditioned Probability (ours)
Refer to caption Tell me a bio of Vincent van Gogh. Refer to caption Vincent van Gogh was a Dutch post-Impressionist painter who is widely considered one of the most important figures in the history of art. He was born on March 30, 1853, in Zundert, Netherlands, and died on July 29, 1890, in Auvers-sur-Oise, France. Van Gogh’s early life was marked by a series of unsuccessful careers, including as a teacher, a pastor, and an art dealer. He began to focus on art in his late 20s, and in 1886, he moved to Paris, where he became involved with the avant-garde art scene. In Paris, van Gogh met many other artists, including Claude Monet, Pierre-Auguste Renoir, and Paul Gauguin, who would become his closest friend and mentor . Gauguin’s influence led van Gogh to adopt a more vibrant and expressive style, characterized by bold colors and thick brushstrokes.
Figure 11: Claim-Conditioned Probability vs Maximum Probability baseline.
Maximum Probability
Refer to caption Tell me a bio of Bill Gates. Refer to caption Bill Gates, born William Henry Gates III on October 28, 1955, is an American entrepreneur, software developer, and philanthropist. He is best known as the co-founder of Microsoft Corporation, one of the world’s largest and most successful software companies. His net worth is estimated to be around \$131.7 billion and is consistently ranked among the world’s wealthiest people. Gates grew up in Seattle, Washington , with his parents, an attorney and a schoolteacher , and two sisters . He showed an early aptitude for mathematics and computer programming and attended Lakeside School, a private preparatory school in Seattle where he met and befriended his future business partner, Paul Allen. In 1973, Gates enrolled at Harvard University, but he soon found himself uninterested in his classes and spent much of his time on programming projects. He eventually dropped out of Harvard to focus on developing Microsoft. Gates and Allen started Microsoft in 1975 , initially providing software for the Altair 8800, an early personal computer.
Claim-Conditioned Probability (ours)
Refer to caption Tell me a bio of Bill Gates. Refer to caption Bill Gates, born William Henry Gates III on October 28, 1955, is an American entrepreneur, software developer, and philanthropist. He is best known as the co-founder of Microsoft Corporation, one of the world’s largest and most successful software companies. His net worth is estimated to be around \$131.7 billion and is consistently ranked among the world’s wealthiest people. Gates grew up in Seattle, Washington, with his parents, an attorney and a schoolteacher, and two sisters. He showed an early aptitude for mathematics and computer programming and attended Lakeside School, a private preparatory school in Seattle where he met and befriended his future business partner, Paul Allen. In 1973, Gates enrolled at Harvard University, but he soon found himself uninterested in his classes and spent much of his time on programming projects. He eventually dropped out of Harvard to focus on developing Microsoft. Gates and Allen started Microsoft in 1975, initially providing software for the Altair 8800, an early personal computer.
Figure 12: Claim-Conditioned Probability vs Maximum Probability baseline.
Appendix H Resources and Expenses
For a single run of data generation and UQ methods evaluation, we spent 12 days of Nvidia A100 GPU compute time. OpenAI API was used mainly for splitting and matching atomic claims, the total cost for all English models (Mistral 7b, Vicuna 13b, Jais 13b, GPT-3.5-turbo) was 40\$. The cost of generating English biographies with GPT-3.5-turbo was 13\$.